OpenStudy (anonymous):

\[\huge if\quad \frac { x-1 }{ x } =\quad y\quad ,\quad and\quad \frac { y+1 }{ y } =\quad x\quad \\ then\quad find\quad the\quad value\quad of\quad 'x-y'\quad IS\quad :\]

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

U mean x-y

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

\[\[\huge if\quad \frac { x-1 }{ x } =\quad y\quad ,\quad and\quad \frac { y+1 }{ y } =\quad x\quad \\ then\quad find\quad the\quad value\quad of\quad 'x-y'\quad IS\quad :\]\]

OpenStudy (anonymous):

Unable to read....... Can u PLZ draw it

OpenStudy (anonymous):

top of u

OpenStudy (anonymous):

MATH PROCESSING ERROR

OpenStudy (anonymous):

From first u will get xy=x-1 From second u will get xy=y+1 THUS, x-1=y+1 x-y=2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome